x to the 2 power- 3 - 3/4 = 0
a) solve by completing the square
b) solve using the quadratic formula
(x to the 2 power minus x minus three-fourths equal 0.

Answer by stanbon(75887)

x to the 2 power- 3x - 3/4 = 0
---
a) solve by completing the square
x^2 - 3x - 3/4 = 0
x^2 - 3x + ? = (3/4)+?
x^2 - 3x + (3/2)^2 = (3/4) + (3/2)^2
(x-(3/2))^2 = (3/4) + (9/4)
(x-(3/2))^2 = 12/4=3
x-3/2 = sqrt(3) or x-3/2 = -sqrt(3)
x = (3/2)+sqrt(3) or x = (3/2)-sqrt(3)
---------------------------------------------------------
b) solve using the quadratic formula
a = 1; b = -3; c = -3/4
---
x = [-b +- sqrt(b^2-4ac)]/(2a)
x = [3 +- sqrt(9 - 4*(-3/4)]/2
x = [3 +- sqrt(12)]/2
x = (3/2) + (2/2)sqrt(3) or x = (3/2)-(2/2)sqrt(3)
x = (3/2)+ sqrt(3) or x = (3/2)-sqrt(3)
============================================
Cheers,
Stan H.